a radical sign with 12x + 12(under it) + a radical sign with 3x + 3 ( under it)
Answer by ilana(307)

I am going to write your problem as sqrt(12x+12)+sqrt(3x+3).
Now, 12x+12=4(3x+3), and since sqrt(ab)=sqrt(a)*sqrt(b),
sqrt(12x+12)=sqrt(4(3x+3))=sqrt(4)*sqrt(3x+3)=2sqrt(3x+3).
So we want 2sqrt(3x+3)+sqrt(3x+3)=2*sqrt(3x+3)+1*sqrt(3x+3)=3sqrt(3x+3).